OpenStudy (janu16):

The lateral area of a cube is 36 square inches. How long is each edge?

OpenStudy (aaronq):

\(Area_{cube}=side^2=36~in^2\) \(side=\sqrt{36~in^2}\)

OpenStudy (janu16):

So what do you do 36 divide by 3?

OpenStudy (janu16):

I mean 36/4?

OpenStudy (aaronq):

the area is given by the product of two sides of a square, or equivalently, the square of the side. \(A=side*side=side^2\) if we know the area is 36 in^2, we can take the squared root of this number to get the length of one side.

OpenStudy (janu16):

So is it gonna be 3?

OpenStudy (aaronq):

no, what's the squared root of 36? \(\sf \huge side =\sqrt{36}\)

OpenStudy (aaronq):

in other words, whats a number when multiplied by itself gives you 36?

OpenStudy (janu16):

6 but thats not my answer choice

OpenStudy (aaronq):

that's the answer though..

OpenStudy (janu16):

I think u do square root of 6

OpenStudy (aaronq):

oh it's a cube, i'm sorry i didn't read the question right.

OpenStudy (aaronq):

so it's all sides of a cube, there are 6 sides on a cubes surface. so first we have to divide 36 by 6 which gives 6 then we take the squared root

OpenStudy (janu16):

Its fine and thanks i will give you a medal

OpenStudy (aaronq):

you are right, squared root of 6 shoujld be the answer

OpenStudy (janu16):

You do 36 divided by 4 than square root 9 to get 3